7. The position of a particle is given by the function s(t) = 100 - 151-4.91. Estimate the particles instantaneous rate of
Answer #1

Method 1 Position S(t) = 100-150- 4.96 Instantaneous rate of change of distance ds d (100-15t - 4.96²) dt dt = -15-2(4.9+) =